HUBBLE SPACE TELESCOPE DAILY REPORT #1896 PERIOD COVERED: 0000Z (UTC) 06/10/97 - 0000Z (UTC) 06/11/97 Daily Status Report as of 162/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6887 (A Photometric Search for Faint Companions to Nearby Stars) The WF/PC-2 was used to make a photometric search observations of the nearby star ROSS248 in order to look for possible companions. The technique involves use of the Planetary Camera with wide band pass filters to photometrically measure for the presence of resolved dark companions. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Seven Sets of STIS/CCD 7675 (STIS Non-Scripted Parallel Archive Proposal) The Space Telescope Imaging Spectrograph (CCD) was used in the parallel mode to obtain slitless spectroscopy with the G750L grism to search for emission line galaxies at intermediate to high redshift. The observations were completed as planned, and no anomalies were reported. 1.3 Completed Two Sets of NIC/1/2/3 7710 (NICMOS MULTIACCUM Dark Characterization 8/8) The purpose of this proposal is to obtain group 8 of the on-orbit darks for the MULTIACCUM predefined sequences requested by GOs, before the start of Cycle 7. The observations completed with no reported problems. 1.4 Completed Three Sets of NIC/1/2/3 7703 (NICMOS MULTIACCUM Dark Characterization 1/8) The purpose of this proposal is to obtain group 1 of the on-orbit darks for the MULTIACCUM predefined sequences requested by GOs, before the start of Cycle 7. The observations completed with no reported problems. 1.5 Completed Two Sets of STIS/MA1 7093 (Dark Measurement for MAMA Band 1, and Band 2) The Space Telescope Imaging Spectrograph (MA1) was used to make dark calibration observations of the MAMA Bands 1 and 2. These observations are to be used to establish that the detectors are performing nominally, that no internal movement of the MCP has taken place, and to establish the dark level in orbit for a given operating temperature. The observations were completed as planned, and no anomalies were reported. 1.6 Completed WF/PC-2 6194 (WF/PC-2 Cycle 5: Polarization Proposal) The WF/PC-2 was used to make polarization observations of star 09-001 at four different spacecraft role angles. This should permit a detailed and independent calibration of each polarizer quad. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Two Sets of NIC/1/2/3 7038 (NICMOS Target Acquisition Test) This test will verify the ability of NICMOS and NSSC-1 flight software to perform Mode-2 (onboard autonomous) target acquisitions in support of coronographic imaging. The efficacy of target location, image centrartion and coordinate transformation (field distortion an SI-to-FGS) on-board software will be evaluated. Flight software data table updates derived from the pre- and post-target acquisition images and FGS data may follow the execution of this test. The test completed nominally with no reported problems. 1.8 Completed Two Sets of NIC/1/2/3 7704 (NICMOS MULTIACCUM Dark Characterization 2/8) The purpose of this proposal is to obtain group 2 of the on-orbit darks for the MULTIACCUM predefined sequences requested by GOs, before the start of Cycle 7. The observations completed with no reported problems. 1.9 Completed NIC/2 7330 (The Fueling of Active Nucleii: A NICMOS Snapshot Survey) The Near Infrared Camera (Camera Two) was used to probe the dark matter distribution of M31 far from the center by observing intervening absorption from the MgII doublet (lambda 2800) in the spectrum of one background QSO located in a geometrically optimal position. HST observations are required because this feature is in a wavelength range not accessible from the ground, and the QSO magnitudes and expected strength of absorption make IUE observations unfeasible. The proposal completed as expected with no anomalies. 1.10 Completed WF/PC-2 7553 (Pure Parallel Archive Survey II) WF/PC-2 parallel observations were obtained in the ultraviolet in search for interesting objects. This program is designed to find extragalactic sources with near ultraviolet excesses (3000 A0) outside the galactic plane and far ultraviolet sources (1600 A0) in the galactic plane. The procedure was executed as scheduled, and no anomalies were reported. 1.11 Completed WF/PC-2 6619 (High Resolution Images of QSO Lyman Alpha Absorbing Galaxies) The WF/PC-2 was used to make several images of a galaxy field (Field 1139) in order to obtain high resolution observations of the Lyman Alpha absorbing galaxies within this field-of-view. The purpose is to directly identify the objects responsible for low and intermediate redshift QSO Lyman-Alpha systems. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed Five Sets of NIC/1/2/3 7701 (NICMOS Short One Orbit Parallel) The Near Infrared Camera (Cameras One, Two and Three) was used as part of a short NICMOS one-orbit parallel program, to obtain "JHK" imaging of sky regions when another instrument is prime. The observations were completed as scheduled, and no anomalies were reported. 1.13 Completed WF/PC-2 6553 (Dissecting 3C-324: Anatomy of an Aligned Radio Galaxy at Z=1.206) The WF/PC-2 was used to make observations of the aligned radio galaxy 3C324 which is known to have a red shift of Z=1.206. This is a prototypical example of the radio galaxy "alignment effect", wherein the spatial orientation of the rest-frame of the ultraviolet continuum emission is closely aligned with the major axis of the FRII radio source. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed Four Sets of STIS/CCD 7700 (STIS Non-Scripted Parallel Archive Proposal) The Space Telescope Imaging Telescope (CCD) was used to obtain slitless spectroscopy with the G750L grism to search for emission line galaxies at intermediate to high redshift. Field galaxy surveys have shown that at high redshift, most galaxies have strong emission lines, and are compact and blue. This suggests that star formation rates were higher than at present,and that AGNs may have been more frequent. The low background and high resolution of HST separates these objects from the background. The proposal completed as planned, and no problems were reported. 1.15 Completed Four Sets of WF/PC-2 6909 (WF/PC-2 Cycle 6 Calibration: Earth Flats The WF/PC-2 was used to make observations of the bright Earth to obtain streak flats which are used to construct high quality flat fields for the WF/PC-2 filter set. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed Three Sets of NIC/1/2/3 7676 (NICMOS-One Orbit Parallel) The Near Infrared Camera (Cameras One, Two and Three) was used to conduct an on-orbit calibration procedure. This was done in parallel with the prime science observations. The procedure was completed as scheduled, and no anomalies were noted. 1.17 Completed STIS/CCD 7063 (STIS SMOV CCD Contamination Monitor) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of standard star GD153 in order to measure flux and look for any gross changes over time that might indicate contamination. The observations were completed as scheduled, and no anomalies were reported. 1.18 Completed Two Sets of WF/PC-2 6802 (Cosmology With the Medium Deep Survey) The WF/PC-2 was used to make observations taken in parallel with the prime science observations as part of the Medium Deep Sky Survey. The observations were executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 11 Successful : 11 2.3 Operations Notes: Battery #4 discharge was started at 161/21:14Z. The process is proceeding nominally. The discharge is expected to take approximately 15 to 18 hours. The NASCOM MDM OTUs were turned off at 162/00:38Z affecting events during a period through 161/01:28Z. All data was recovered from the SSR. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science and SMOV operations. /DMH